### v1

# Code Clone Detection via an AlphaFold-Inspired Framework

Abstract.
Code clone detection, which aims to identify functionally equivalent code fragments, plays a critical role in software maintenance and vulnerability analysis. Substantial methods have been proposed to detect code clones, but they fall short in capturing code semantics or relying on language-specific analyzers. Inspired by the remarkable success of AlphaFold in predicting three-dimensional protein structures from protein sequences, in this paper, we leverage AlphaFold for code clone detection based on the insight that protein sequences and token sequences share a common linear sequential structure. In particular, we propose AlphaCC, which represents code fragments as token sequences to ensure multi-language applicability and adapts AlphaFold’s sequence-to-structure modeling capability to infer code semantics. The pipeline of AlphaCC goes through three steps. First, AlphaCC transforms each input code fragment into a token sequence and, motivated by AlphaFold’s use of multiple sequence alignment (MSA) to enhance contextual understanding, constructs an MSA from lexically similar token sequences. Second, AlphaCC adopts a modified attention-based encoder based on AlphaFold to model dependencies within and across token sequences. Finally, unlike AlphaFold’s protein structure prediction task, AlphaCC computes similarity scores between token sequences through a late interaction strategy and performs binary classification to determine code clone pairs. Comprehensive evaluations on three language-diverse datasets demonstrate AlphaCC’s applicability across multiple programming languages. On two semantic clone detection datasets, it consistently outperforms all baselines, showing strong semantic understanding. Moreover, AlphaCC maintains competitive efficiency, enabling practical usage in large-scale clone detection tasks.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Software Engineering; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/06copyright: none
1. INTRODUCTION
Code clones, which refer to code fragments that are identical or equivalent in functionality, are frequently introduced through copy-and-paste practices or repeated implementation of similar program logic (Baker, 1995; Kim et al., 2005) in software engineering. While such reuse can reduce engineering effort and accelerate software development, it may also introduce hidden maintenance risks. Such risks materialize when duplicated code originates from buggy source code, as undetected defects propagate across the codebase, leading to accumulating technical debt and increasing maintenance cost (Kim et al., 2017). Furthermore, many buggy clones are not exact copies but modified versions of the original code, making the underlying bugs harder to detect and eliminate (Li and Ernst, 2012). To this end, code clone detection is essential for identifying and addressing issues that may exist in semantically equivalent code fragments distributed throughout the codebase.
Various approaches have been proposed for code clone detection. Based on the intermediate representation used, these methods can be divided into four main categories (Jiang et al., 2007; White et al., 2016): text-based, token-based, Abstract Syntax Tree (AST)-based, and graph-based methods. The latter two categories of methods (AST‑based and graph‑based) extract structured intermediate representations (Jiang et al., 2007; Zhang et al., 2019; Xu et al., 2024; Yu et al., 2023; Zhao and Huang, 2018; Hua et al., 2020; Wu et al., 2020), such as abstract syntax trees and program dependency graphs, which are well-suited for detecting semantic clones (Wang et al., 2023), namely code fragments that perform the same or very similar functionality but differ syntactically (Yu et al., 2019). However, both methods rely on language-specific analyzers for structural extraction and suffer from low efficiency. The lack of corresponding analysis tools to support reliable construction of ASTs or program graphs is especially serious for some domain-specific languages (DSLs), such as SageMath (sag, 2025). Considering scenarios such as hardware design and data analytics, where the use of DSLs is steadily increasing (Damyanov and Sukalinska, 2015), AST-based and graph-based methods become impractical. In contrast, text-based and token-based methods employ relatively simple intermediate representations: the former treats source code as plain text, while the latter performs lexical analysis to transform source code into a sequence of tokens (Baker, 1997; Ankali and Parthiban, 2021; Feng et al., 2024). These lightweight representations offer advantages in efficiency and broad language applicability, but sacrifice the capability to capture code semantics (Sajnani et al., 2016; Kamiya et al., 2002). In summary, existing clone detection methods suffer from either limited capability for semantic clone detection or strong reliance on language-specific analyzers.
To address these limitations, we aim to develop a new code clone detection framework that features both semantic understanding capability and multi-language applicability. Our inspiration comes from AlphaFold (Jumper et al., 2021), which has achieved remarkable success in the biological domain by accurately predicting three-dimensional protein structures from protein sequences. The success of AlphaFold demonstrates its ability to abstract rich structural representations from linear symbolic sequences by capturing complex sequence dependencies that govern functional organization (Chen et al., 2024; Wang et al., 2024). Motivated by this, we explore whether AlphaFold can be applied to code clone detection, based on the fact that code fragments can likewise be represented as symbolic sequences, namely token sequences. Interestingly, token sequences and protein sequences share several key similarities:
  • Linear symbolic structure: Both are linear sequences of discrete symbols, i.e. tokens and amino acids, respectively.
  • Inner-sequence interaction: The meaning of each token or amino acid in a sequence depends on its sequence neighborhood.
  • Cross-sequence interaction: Homologous protein sequences are analogous to lexically similar token sequences, which provide context for learning underlying function.
Building on these similarities, we introduce a new framework to detect code clones that, on the one hand, benefits from a token-based representation without dependence on language-specific analyzers, and on the other hand, leverages the sequence-to-structure modeling capability of AlphaFold to capture code semantics.
For an input protein sequence, AlphaFold first constructs a multiple sequence alignment (MSA) to incorporate information from homologous protein sequences, thereby enriching the input representation with co-evolutionary features (Rao et al., 2021). Then, an attention-based encoder, Evoformer, is employed to model complex dependencies, enabling information exchange both within individual sequences and across homologous sequences (Hu et al., 2022). The protein modeling process in AlphaFold inspires the design of a code clone detection framework that combines MSA construction and an attention-based encoder to enrich intermediate token representations and improve the capability to detect semantic clones.
In this paper, we propose AlphaCC, an AlphaFold-inspired framework for code clone detection. Given a pair of code fragments, AlphaCC begins by preprocessing them into token sequences. It then proceeds through three main stages: a token semantic enhancer, an attention-based encoder, and a classification and optimization module. The token semantic enhancer aims to enrich token representation of code fragments. First, inspired by the MSA construction of AlphaFold, for each input token sequence, several lexically similar token sequences are retrieved from large-scale codebase together with their code context to construct a multiple sequence alignment of code (Code MSA). Then, AlphaCC introduces a novel mechanism that projects each token category into a type-specific semantic space, enabling the extraction of fine-grained, type-aware semantic features. Finally, an additional attention mechanism is applied to integrate information across these semantic spaces, enabling interactions between different token types. Building on these enriched token embeddings, an attention-based encoder adapted from AlphaFold’s Evoformer, termed Codeformer, is designed to integrate inner-sequence attention and cross-sequence attention, thereby capturing dependencies both within individual token sequence and across multiple token sequences. At the final stage, the classification and optimization module computes fine-grained similarity scores between code fragments via late interaction and performs binary classification to determine whether a given pair constitutes a code clone. To supervise this process, a margin-based loss function is employed as the optimization objective, increasing scores for semantically similar code fragment pairs while decreasing scores for dissimilar ones. It should be emphasized that the final stage is specific to the code clone detection task and is irrelevant to AlphaFold.
To evaluate the effectiveness, multi-language applicability, and efficiency of AlphaCC, we conduct experiments on three widely used benchmarks for code clone detection: GCJ (gcj, 2017), BigCloneBench (big, 2020), and OJClone (Mou et al., 2016). Among them, GCJ and BigCloneBench are datasets based on Java, whereas OJClone consists of code written in C. The experimental results demonstrate that AlphaCC consistently outperforms all baseline methods across the three benchmarks, achieving clone detection accuracies of $97.4\%$, $96.0\%$, and $98.6\%$ on GCJ, BigCloneBench, and OJClone, respectively. Moreover, AlphaCC exhibits superior efficiency, achieving faster detection than all AST-based and graph-based approaches under identical computational settings.
Overall, our contributions are as follows:
  • We observe several key similarities between protein sequences and token sequences, and creatively adapt the core idea of AlphaFold, namely sequence-to-structure modeling, to the task of code clone detection.
  • We propose AlphaCC, a code clone detection framework that adopts token representation and, inspired by AlphaFold, incorporates multiple sequence alignment construction and a modified attention-based encoder.
  • We conduct comprehensive experiments on three widely used benchmarks, demonstrating the effectiveness, multi-language applicability, and efficiency of AlphaCC.
The remainder of the paper is organized as follows. Section 2 reviews related work on code clone detection. Section 3 introduces the biological inspiration behind our method. Section 4 presents AlphaCC, an AlphaFold-inspired framework for code clone detection. Section 5 details the experimental setup and results to evaluate the effectiveness, multi-language applicability, and efficiency of AlphaCC. Section 6 discusses the possible interpretations of AlphaCC and examines potential threats to validity. Section 7 concludes this paper.
Figure 1. From Protein to Code
Figure 2. From AlphaFold to AlphaCC
2. RELATED WORK
This section introduces related works on code clone detection. According to different intermediate representations of code, existing code clone detection methods can be divided into four categories: text-based, token-based, AST-based, and graph-based.
Text-based methods treat source code as plain text and employ string matching techniques to identify clones. Johnson (Johnson, 1993) designs a prototype system based on the fingerprints of code to identify code clones. Baker proposes Dup (Baker, 1995), which detects duplicate code by converting code strings into parameterized suffix trees and then identifying matches within these trees. Despite their rapid detection speed, the absence of program syntax weakens text-based methods in handling complex syntactic or semantic clones.
Token-based methods typically transform source code into code tokens by lexical analysis first, then detect code clones through the comparison of tokens. Kamiya et al. (Kamiya et al., 2002) develop CCFinder to match code snippets token-by-token using suffix-tree matching algorithm. Sajnani et al. (Sajnani et al., 2016) implement SourcererCC to detect code clones by building an optimized inverted index for code tokens. While token-based methods are highly efficient, their lack of syntactic information significantly limits their effectiveness in detecting semantic clones.
AST-based methods use the abstract syntax trees of code snippets to detect code clones. Baxter et al. (Baxter et al., 1998) split the original AST into subtrees and compare the hash code of subtrees to detect code clones. Jiang et al. (Jiang et al., 2007) design DECKARD to compute characteristic vectors for AST subtrees and employ locality-sensitive hashing to group similar vectors, thereby identifying clones. Furthermore, some methods, such as CDLH (Wei and Li, 2017), ASTNN (Zhang et al., 2019), and DSFM(Xu et al., 2024), use deep learning models to encode AST into vectors and then calculate the similarity between vectors to detect code clones. While AST-based approaches can identify semantic code clones, they are often hindered by efficiency issues and limited multi-language applicability because of heavy reliance on language-specific AST analyzers.
For graph-based methods, they first analyze different graph representations, such as control flow graph (CFG) and program dependency graph (PDG), for code, and then further process the representation data to detect code clones. Krinke (Krinke, 2001) finds similar subgraphs in PDG to detect code clones. Wu et al. (Wu et al., 2020) design SCDetector, which uses centrality analysis to detect code clones from CFG. Similar to AST-based methods, graph-based approaches are effective in capturing semantic similarities. However, graph-based approaches face low detection efficiency, and their dependence on language-specific graph construction tools further restricts widespread adoption.
In summary, existing clone detection methods are constrained by insufficient ability of semantic clone detection or strong dependence on language-specific analyzers. These limitations motivate the design of our new framework, AlphaCC.
3. Biological Inspiration: From Protein Structure Prediction to Code Clone Detection
This section introduces the biological inspiration behind our approach. Section 3.1 draws an analogy between protein sequences and token sequences, while Section 3.2 compares the framework of AlphaFold and AlphaCC.
3.1. From Protein to Code: Anology of Sequence
We establish an analogy between protein sequences and token sequences from three key perspectives.
First, as shown in Figure 1, protein sequences and token sequences share a common structural pattern: a linear symbolic structure. That is, both protein sequences and token sequences are linear sequences consisting of a finite number of discrete symbols—amino acids in the former and code tokens in the latter. Moreover, there are $20$ standard types of amino acids in protein sequences, such as Alanine, Cysteine, Lysine, and Proline. Similarly, for a given programming language, code tokens fall into a finite number of categories, such as modifiers, identifiers, keywords, and operators. Each token belongs to exactly one category, as determined by the language’s lexical analyzer.
Second, both protein sequences and token sequences exhibit inner-sequence interaction, where the meaning of each element is context-dependent and determined by its neighboring elements within the sequence. In a protein sequence, the role of an amino acid is dependent on its neighboring residues. For instance, in the lysozyme protein sequence, the segment from positions $25$ to $35$ forms an $\alpha$-helix structure and consists of the following amino acids:
L25-G26-L27-W28-V29-A30-W31-R32-N33-R34-F35,
where every letter denotes an amino acid and the accompanying number specifies its position. The relative positioning of amino acids such as W28 and W31 contributes to the formation of a hydrophobic core that stabilizes the $\alpha$-helix structure, while the local arrangement of W28–V29–A30 creates a hydrophobic patch that is essential for local stability and contributes to the protein’s overall function. Similarly, in a token sequence, the semantic role of each token is determined by its surrounding token context within the sequence, as program semantics often arise from token patterns rather than individual tokens. For example, considering the C‑style loop
for(int a = 0; a < N; a++),
its corresponding token sequence is
["for", "(", "int", "a", "=", "0", ";",
[…]
In the token sequence, token "a" appears multiple times, and its interpretation—whether as a declaration, a condition variable, or an increment target—relies on the surrounding tokens such as "int", "<", and "++".
Finally, both protein sequences and token sequences exhibit cross-sequence interaction, where related sequences provide mutual context for understanding the structure of proteins and the semantics of code, respectively. For protein sequences, homologous sequences exhibit similar three-dimensional structures. A well-known example involves hemoglobin and myoglobin, whose protein sequences are homologous and belong to the globin family. They share a three-dimensional structure, known as the globin fold, which consists of eight $\alpha$-helices arranged in a compact, globular configuration. While lexically similar token sequences may differ in variable names, statement ordering, or syntactic constructs, they can still implement the same underlying computational logic and yield identical functional behavior.
3.2. From AlphaFold to AlphaCC: Anology of Framework
We draw an analogy between the frameworks of AlphaFold and AlphaCC to emphasize their connection and distinction.
AlphaFold is a representative end-to-end framework that takes a protein sequence as input and predicts its three-dimensional structure. As illustrated in Figure 2, the overall framework consists of three stages: MSA construction for evolutionary relationships modeling, the Evoformer module for feature extraction, and the structure module for final three-dimensional structure prediction. Since the structure module in AlphaFold serves as a decoder that maps learned features to three-dimensional coordinates—a stage irrelevant to code clone detection—AlphaCC focuses on adapting the first two stages to extract semantic features from token sequences.
In the MSA construction stage, AlphaFold builds a multiple sequence alignment for each protein sequence to compensate for the limited evolutionary information available from a single sequence (Zhang et al., 2024b). AlphaCC adopts a similar design, constructing a multiple sequence alignment of code for each token sequence to enhance semantic understanding. To further enrich token representation, AlphaCC additionally introduces type-specific projections and an interaction mechanism.
Based on MSA, AlphaFold employs Evoformer to capture two crucial types of information: positional information of amino acids within each sequence and interactions among aligned amino acids across the entire MSA (Lupo et al., 2022). Inspired by Evoformer, AlphaCC introduces inner-sequence self-attention and inter-sequence cross-attention to process Code MSAs. These two attention modules capture structural dependencies within individual token sequences, such as control and data flows, and identify co-occurrence patterns across aligned code tokens.
In the final stage, the Structure module from AlphaFold is removed. Instead, to tailor AlphaCC for code clone detection, the token-level similarity between code fragments is calculated through late interaction mechanism.
4. AlphaCC
Building on cross-domain adaptations from AlphaFold, the framework of AlphaCC enables the detection of semantically equivalent code pairs across syntactic and structural variations, with broad applicability across programming languages. To support this, our method is implemented through a three-stage process, which is described in the following sections.
4.1. Token Semantic Enhancer
Raw token sequences lack the structural information necessary for deep semantic understanding. To address this, we design the Token Semantic Enhancer, which enriches token sequences into semantically enhanced representations through multiple sequence alignment of code (Code MSA), enabling the model to better capture the underlying intent and functionality of the code.
Codebase Preparation. To construct Code MSAs for token sequences, a large-scale and high-quality codebase is required. As our experimental datasets (see Section 5) involve programs written in Java and C, we build separate codebases for each language. Specifically, we collect their top 10,000 GitHub repositories, ranked by the criticality score, which reflects a project’s overall impact and importance (cri, 2024; Dou et al., 2024). Since the target datasets are function-level, we parse all source code from these repositories into individual functions. This results in a total of 68,093,823 functions for Java and 129,219,483 functions for C.
To obtain vector representations of tokens, we employ Word2Vec (Mikolov et al., 2013), a widely adopted and computationally efficient embedding technique. Compared to large language models, Word2Vec offers a lightweight and practical alternative for unsupervised token representation learning. All source code, including that from the experimental datasets and the constructed codebases, is first tokenized using javalang(jav, 2022) for Java and pycparser(pyc, 2024) for C. The resulting tokens are then used to train the Word2Vec model, which generates embedding vectors for individual tokens.
Figure 3. Construction of Code MSA
Figure 4. Pipeline of Encoding Code MSA in AlphaCC
Construction of Code MSA. Inspired by AlphaFold’s use of Multiple Sequence Alignment (MSA) to reveal conserved evolutionary patterns in proteins, we introduce multiple sequence alignment of code (Code MSA) to uncover shared structural patterns across syntactically diverse code implementations. As shown in Figure 3, for each input code fragment, we employ an efficient $N$-gram-based similarity search to retrieve four lexically similar token sequences from a candidate group. $N$-gram is adopted for its simplicity and efficiency in capturing recurring token subsequences, a key feature of code clones (Smith and Horwitz, 2009; Martinez-Gil, 2024; Nakagawa et al., 2021). Each sequence is represented using $5$-grams and compared via cosine similarity. To ensure uniformity, all sequences are standardized to a fixed length. For shorter codebase sequences, we incorporate surrounding code as contextual padding. Prior work has demonstrated that such context provides crucial information for understanding the purpose and behavior of a code fragment (Cheng et al., 2022; Zhang et al., 2024a; Bansal et al., 2021). All retrieved token sequences, together with the original sequence, constitute a Code MSA.
This procedure enriches the model’s input by expanding it from a single implementation to a diverse set of representative patterns. For instance, when comparing recursive and iterative versions of a greatest common divisor (GCD) function, Code MSA provides multiple structurally similar examples from both approaches. By aligning these variations, the model learns to abstract the core computational logic, such as modulo operations and base case conditions. Simultaneously, it ignores superficial syntactic variations like variable names or control constructs. This enables a more robust understanding of functional semantics.
Table 1. $15$ token types
"Separator", "Identifier", "Operator", "Keyword",
"Modifier", "DecimalInteger", "BasicType", "String",
"Boolean", "Null", "DecimalFloatingPoint",
"Annotation", "HexInteger", "HexFloatingPoint", "OtherType"
Type-Aware Mechanism for Feature Fusion. To further enhance the model’s structural awareness, we introduce a type-aware mechanism. As summarized in Table 1, following previous research (Dou et al., 2024), all tokens are first categorized into 14 lexical types (e.g., Identifier, Operator, Keyword), along with an OtherType that accounts for rare tokens comprising less than 1% of the dataset. Each type is assigned a unique identifier. Based on the standardized Code MSAs, we extract the corresponding token type IDs and transform them into continuous vectors via an embedding layer. In parallel, the tokens themselves are encoded using a pretrained Word2Vec model. The resulting token and type embeddings are combined through element-wise addition, yielding a unified Code MSA matrix that captures both token semantic and structural information.
To model type-specific semantics, we apply a type-specific projection layer, where tokens belonging to the same lexical type are projected into a shared semantic subspace. This encourages the model to learn type-consistent representations, for example, all Operator tokens are embedded within the same subspace, enabling the model to capture their common functional behavior. Finally, a token-type aware attention mechanism is employed to integrate information across these subspaces, effectively modeling interactions and dependencies between different token types.
4.2. Codeformer
To interpret the enriched feature matrix, we introduce Codeformer, an encoder architecture inspired by AlphaFold’s Evoformer. As illustrated in Figure 4, Codeformer employs a dual-attention mechanism that models dependencies along two aspects:
Inner-Sequence Self-Attention. This component operates row-wise across the Code MSA matrix, where each row represents a single token sequence. It captures intra-procedural logic, including control flow (e.g., if-else branches) and data flow (e.g., variable definitions and usages). Instead of computing separate attention maps for each sequence, Codeformer aggregates them into a shared attention map applied across all sequences. This design leverages the structural homology often present among code fragments (as discussed in Section 3.1), allowing the model to identify consistent interaction patterns that would be overlooked by isolated self-attention.
Inter-Sequence Cross-Attention. This layer operates column-wise to enable interactions among tokens at aligned positions across different code fragments. It facilitates the identification of semantic equivalence across syntactic variations. For example, the model can align a for loop in one sequence with a semantically equivalent while loop in another, by matching their initialization, condition, and update components, even when expressed differently. This mechanism is essential for recognizing semantically similar code that has undergone significant refactoring.
Through the coordinated application of these two attention mechanisms, Codeformer generates token representations that are sensitive to both the internal logic of individual procedures and the external variation across functionally equivalent implementations.
4.3. Classification and Optimization
The final stage is responsible for comparing the encoded representations and optimizing the model for the clone detection task.
Figure 5. Late Interaction
Late Interaction. Previous similarity computation methods typically compress entire code fragments into single vectors before comparison, a process that often results in significant information loss. To address this limitation, as shown in Figure 5, AlphaCC adopts a Late Interaction strategy (Feng et al., 2020; Khattab and Zaharia, 2020; Santhanam et al., 2021). Instead of collapsing sequence information early, we retain the token-level representations generated by Codeformer and compute similarity scores through fine-grained, token-wise comparisons.
Specifically, we first average the token embeddings at each aligned position across all sequences in the Code MSA to obtain a single pooled representation. All token vectors are then normalized to ensure consistent scaling. Let $\mathcal{C}_{1}=\{\mathbf{t}_{1}^{(1)},\mathbf{t}_{2}^{(1)},\dots,\mathbf{t}_{n}^{(1)}\}$ and $\mathcal{C}_{2}=\{\mathbf{t}_{1}^{(2)},\mathbf{t}_{2}^{(2)},\dots,\mathbf{t}_{n}^{(2)}\}$ denote the sets of normalized token embeddings for two code fragments $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively, where each $\mathbf{t}_{i}^{(1)},\mathbf{t}_{j}^{(2)}\in\mathbb{R}^{d}$ and $d$ is the embedding dimension. The Euclidean distance between two token vectors is defined as:
(1) $\text{dist}(\mathbf{t}_{i}^{(1)},\mathbf{t}_{j}^{(2)})=\left\|\mathbf{t}_{i}^{(1)}-\mathbf{t}_{j}^{(2)}\right\|_{2}$
The overall similarity between two code fragments is computed as the mean of the minimum pairwise distances, where each token in $\mathcal{C}_{1}$ is matched to its nearest neighbor in $\mathcal{C}_{2}$:
(2) $\text{sim}(\mathcal{C}_{1},\mathcal{C}_{2})=\frac{1}{n}\sum_{i=1}^{n}\min_{1\leq j\leq n}\text{dist}(\mathbf{t}_{i}^{(1)},\mathbf{t}_{j}^{(2)})$
This late interaction approach enables fine-grained semantic alignment by directly comparing token-level representations, allowing the model to accurately identify correspondences between semantically related tokens across syntactically diverse code fragments.
Margin Loss. To enable high-accuracy clone detection, we adopt Margin Loss (You et al., 2020; Zhan et al., 2025), which encourages a more discriminative decision boundary than standard binary cross-entropy, improving generalization and reducing sensitivity to ambiguous cases.
Given the Late Interaction similarity score $s_{ij}\in[0,1]$ between a pair of code fragments $(\mathcal{C}_{i},\mathcal{C}_{j})$, and a ground-truth label $y_{ij}\in\{-1,+1\}$ indicating whether the pair constitutes a clone, the loss is defined as:
(3) $\mathcal{L}_{\text{margin}}=\max\left(0,\gamma-y_{ij}\cdot(1-s_{ij})\right)$
where $\gamma>0$ is a predefined margin hyperparameter. It enforces separation between positive and negative pairs, enhancing clone discrimination while providing a tolerance region that regularizes training and reduces overfitting.
Table 2. Comparison of different methods on three datasets
GCJ BCB OJClone
Type Method Precision Recall F1 Score Precision Recall F1 Score Precision Recall F1 Score
SourcererCC 74.8 4.2 7.9 97.5 0.6 1.2 27.2 25.2 26.2
NIL 41.0 9.7 15.7 86.4 3.0 5.8 18.0 37.8 24.4
Token-based Toma 68.3 29.1 40.8 80.7 20.7 32.9 71.1 34.0 46.0
PNIAT+PDG 93.4 94.9 94.1 92.0 95.9 93.8 - - -
Graph-based PNIAT+CFG 93.8 96.0 94.9 92.4 96.6 94.5 - - -
Tamer 60.7 10.3 17.5 86.0 2.1 4.1 - - -
TBCCD 97.3 92.7 95.0 96.6 93.1 94.8 95.4 96.8 96.1
ASTNN 97.9 93.0 95.4 97.9 92.8 95.3 97.5 98.3 97.9
AST-based DSFM 97.2 96.3 96.8 95.3 96.0 95.7 95.7 99.4 97.5
Ours AlphaCC 97.6 97.2 97.4 97.7 94.2 96.0 99.7 97.6 98.6
Table 3. Comparison of different methods in F1 Score on different clone types of BigCloneBench
Type Method T1 T2 ST3 MT3 T4
Token-based SourcererCC 100 97.1 63.7 6.2 0
NIL 100 97.4 68.3 20.5 0
Toma 100 100 94.3 87.3 28.1
Graph-based PNIAT+PDG 100 100 99.2 98.9 91.3
PNIAT+CFG 100 100 100 99.2 93.9
AST-based Tamer 100 100 98.3 51.2 2.7
TBCCD 100 100 100 98.3 94.5
ASTNN 100 100 100 99.7 95.1
DSFM 100 100 100 100 95.4
Ours AlphaCC 100 100 100 100 95.7
5. Evaluation
This section presents a comprehensive evaluation of AlphaCC to rigorously assess its capabilities. Our evaluation is structured around four central research questions (RQs):
  • RQ1 (Effectiveness): How does AlphaCC’s performance in code clone detection compare against state-of-the-art baseline methods?
  • RQ2 (Efficiency): How efficient is AlphaCC in terms of both training and prediction time, particularly for large-scale tasks?
  • RQ3 (Framework Analysis): What is the impact of AlphaCC’s core components and configurations on its overall performance?
  • RQ4 (Robustness and Generalization): How robust is AlphaCC’s performance when trained and tested on datasets with different characteristics and data distributions?
5.1. Experiment Setup
Datasets. We conducted experiments on three widely-used datasets. These datasets are strategically chosen to cover different programming languages and diverse clone types, ensuring a thorough and comprehensive evaluation.
  • GCJ (Google Code Jam) (gcj, 2017): A Java dataset derived from the Google Code Jam programming competition, comprising $1,664$ Java projects drawn from $12$ different competition problems. In line with prior work (Zhao and Huang, 2018; Xu et al., 2024), two code snippets that deal with the same problem are treated as clones. It has been confirmed that most clone pairs in GCJ are semantic clone (Zhao and Huang, 2018). We randomly select $1,000$ code snippets for training, $332$ for validation, and $332$ for testing, resulting in nearly $600,000$ code pairs overall.
  • BigCloneBench (big, 2020): A massive Java dataset, which is constructed from IJaDataset (Svajlenko et al., 2014) and contains over $6$ million clone pairs and $260,000$ false clone pairs. All the clone pairs are hierarchically classified into different clone types: Type-1 (T1) clones are lexically identical after comment removal and formatting normalization; Type-2 (T2) clones extend T1 by permitting systematic renaming of identifiers and replacement of literals. The remaining clone pairs are further categorized based on their similarity scores at the line and token levels: Strongly Type-3 (ST3, similarity $\in[0.7,1.0)$) and Moderately Type-3 (MT3, similarity $\in[0.5,0.7)$), while Type-3+4 (T4) exhibit semantic similarity with minimal textual overlap (similarity $\in[0.0,0.5)$(Svajlenko et al., 2014). T4 clones correspond to the notion of semantic clones as emphasized in the GCJ dataset. The scale and diversity of BigCloneBench make it well-suited for evaluating models across both syntactic and semantic clone detection challenges. Following prior work (Wei and Li, 2017; Yu et al., 2019; Xu et al., 2024), we randomly select $8,134$ code snippets for training, $500$ for validation, and $500$ for testing. Finally, we get $124,750$ code pairs for validation and test set and $750,500$ for training set, including $0.3\%$ of T1 clones, $0.1\%$ of T2 clones, $0.2\%$ of ST3 clones, $1.0\%$ of MT3 clones, and $98.4\%$ of T4 clones.
  • OJClone (Mou et al., 2016): A C language code clone dataset. OJClone is built from student submissions to $104$ programming problems (Mou et al., 2016). Prior studies (Xu et al., 2024) have reported that code clone pairs from OJClone are mainly semantic clones, since almost all code snippets differ significantly in syntax. This makes OJClone a challenging benchmark for evaluating models’ ability to capture deep semantic similarity. Following established practice (Wei and Li, 2017; Yu et al., 2019; Zhang et al., 2019), we randomly select $6,500$ code snippets for training, $500$ for validation, and $500$ for testing, yielding $1,079,051$ training pairs and $280,875$ validation/test pairs (Xu et al., 2024; Yu et al., 2019).
Baselines. To comprehensively evaluate the effectiveness of AlphaCC, we select some state-of-the-art baselines across three types of approaches: token-based, AST-based, and graph-based.
  1. (1)
    Token-based
    • SourcererCC (Sajnani et al., 2016): A tool that identifies code clones by computing the overlap of indexed token subsequences.
    • NIL (Nakagawa et al., 2021): A clone detection technique that employs $N$-gram token indexing to match the longest common subsequence.
    • Toma (Feng et al., 2024): An approach that computes six similarity calculation methods based on token sequences to form feature vectors, and leverages a machine learning classifier to determine clone pairs.
  2. (2)
    AST-based
    • Tamer (Hu et al., 2023): A tool that splits the original AST into simple subtrees to detect code clones.
    • TBCCD (Yu et al., 2019): An approach that utilizes tree-based convolution over token-enhanced AST to detect semantic code clones.
    • ASTNN (Zhang et al., 2019): A model that converts AST into subtrees and uses bidirectional RNN for semantic code clone detection.
    • DSFM (Xu et al., 2024): A code clone detection approach that uses GRU to capture the relationships between subtrees split from the original AST.
  3. (3)
    Graph-based
    • PNIAT+PDG and PNIAT+CFG (Yu et al., 2023): A technique designed to convert irregular code graph data and uses attention mechanisms to capture critical tokens within each statement in parallel to detect semantic code clones.
Experimental Settings. In our experiments, all baseline methods are configured using default settings, widely adopted hyperparameters, or values tuned on our datasets to ensure fair comparison. For our model, we set the number of training epochs to $1$, the learning rate to $0.0001$, the margin parameter to $0.5$, and the token embedding dimension to $256$. All experiments are conducted on a unified computational platform equipped with a 192-core CPU, 755 GiB of RAM, and an NVIDIA L40S GPU to ensure consistency and fairness in comparison.
Evaluation Metrics. To measure the effectiveness of all the models, we adopt the widely used metrics: precision, recall, and F1 score. Precision is the proportion of correctly predicted clone pairs among all predicted clone pairs, while recall measures the proportion of correctly predicted clone pairs among all actual clone pairs. F1 score is the harmonic mean of precision and recall, providing a balanced evaluation of model performance.
5.2. RQ1: Effectiveness Comparison
To address RQ1 regarding effectiveness, we conduct a comprehensive evaluation of AlphaCC against several state-of-the-art clone detection methods across three datasets: GCJ, BigCloneBench, and OJClone, with the results shown in Table 2. The best results are marked in bold, while the second-best are underlined. Some entries in Table 2 are marked with a dash (–) because those baseline methods could not be reproduced on OJClone. This is because OJClone is a C dataset, which some baselines do not support or were not originally designed to handle.
It can be seen that AlphaCC consistently achieves the highest F1 scores across all three distinct benchmarks, reaching $97.4\%$ on GCJ, $96.0\%$ on BigCloneBench, and an outstanding $98.6\%$ on OJClone. This consistent superiority across datasets with diverse programming languages and clone types highlights the effectiveness of our approach. Moreover, AlphaCC maintains top-tier precision and recall, demonstrating its balanced ability to both detect true clones and avoid false positives.
Semantic Clone Detection Superiority. The primary challenge in modern clone detection is accurately identifying semantic clones, which exhibit functional equivalence despite significant syntactic divergence. On two semantic clone datasets, GCJ and OJClone, AlphaCC demonstrates outstanding performance. To further analyze AlphaCC’s strength, we conduct a fine-grained analysis on the challenging BigCloneBench dataset, which categorizes clone types by their syntactic similarity (Table 3). AlphaCC still demonstrates leading performance across all clone types. Crucially, for the most difficult T4 clones, AlphaCC achieves a remarkable F1 score of 95.7%, significantly surpassing all other methods. This performance gain stems from two core architectural innovations: Code MSA, which enhances semantic representation by aligning lexically similar token sequences to reveal functional commonalities; and the Codeformer module, which jointly captures inner-sequence semantics and inter-sequence commonalities, enabling robust abstraction of underlying program logic.
Inherent Multi-Language Applicability. Furthermore, AlphaCC’s core design offers a crucial advantage: inherent multi-language applicability. As shown in Table 2, methods dependent on language-specific analyzers, such as PNIAT, fail to process the C-based OJClone dataset. In contrast, AlphaCC operates on universal token sequences, converting source code into a standardized token stream that abstracts away language-specific syntax. This language-agnostic approach enables AlphaCC to deliver outstanding performance on OJClone, highlighting its potential for deployment across diverse software ecosystems.
Table 4. Runtime performance of different methods on analyzing one million code pairs
Type Method Prediction Time Training Time
Token-based SourcererCC 19s -
NIL 10s -
Toma 35s 2283s
Graph-based PNIAT+PDG 185s 24349s
PNIAT+CFG 185s 24347s
AST-based Tamer 107s -
TBCCD 5827s 56195s
ASTNN 8542s 72417s
DSFM 475s 3928s
Ours AlphaCC 76s 9355s
5.3. RQ2: Efficiency Analysis
To investigate RQ2, we evaluate AlphaCC’s computational efficiency, a key requirement for practical deployment on large-scale clone detection tasks. We benchmark all methods on one million randomly selected code pairs from BigCloneBench, reporting the average execution time over three runs. Following prior work (Dou et al., 2024), we distinguish between the training phase (required by deep learning-based methods) and the prediction phase, which reflects real-world usage.
Prediction Time. AlphaCC demonstrates outstanding efficiency during prediction. As shown in Table 4, it is the fastest among all deep learning methods and even outperforms traditional approaches such as Tamer. Unlike approaches with complex model designs, AlphaCC adopts a lightweight encoder built around a few attention layers. Compared to traditional methods that are often CPU-bound and difficult to accelerate, AlphaCC is designed to fully leverage GPU hardware for efficient execution. This combination of architectural simplicity and GPU-friendly design enables AlphaCC to achieve efficient and scalable inference for large-scale clone detection tasks.
Training Time. Although methods such as DSFM exhibit shorter training durations, AlphaCC demonstrates a clear efficiency advantage over most deep learning-based approaches. As reported in Table 4, AlphaCC completes training in $9,355$ seconds, significantly outperforming graph-based methods such as PNIAT+PDG ($24,349$ seconds), as well as AST-based methods like TBCCD ($56,195$ seconds) and ASTNN ($72,417$ seconds). This efficiency is attributed to AlphaCC’s lightweight design. Moreover, the training phase constitutes a one-time offline cost. When viewed in the context of AlphaCC’s fast and accurate inference capabilities, this initial training investment is therefore justified by substantial downstream benefits in performance and scalability.
5.4. RQ3: Framework Analysis
To address RQ3, we conduct a series of experiments to analyze and validate the AlphaCC’s core designs. All evaluations in this section are based on the GCJ dataset for consistency.
Table 5. Performance of AlphaCC using different code input formats
Method Precision Recall F1
5-sequence Code MSA (AlphaCC) 97.8 96.3 97.1
3-sequence Code MSA 97.1 96.4 96.7
Single token sequence 98.4 25.9 41.0
Impact of Code MSA. Code MSA constitutes a central innovation in AlphaCC, aiming to enhance semantic representation by aligning lexically similar token sequences. We compare the default five-sequence Code MSA with two alternatives: a three-sequence configuration and a single-sequence baseline. As shown in Table 5, the five-sequence MSA achieves an F1 score of $97.1\%$, outperforming both the three-sequence variant ($96.7\%$) and the single-sequence baseline ($41.0\%$). This substantial improvement of Code MSA compared with the single-sequence baseline demonstrates that exposing the model to multiple semantically related sequences encourages learning of deeper functional invariants while mitigating sensitivity to syntactic variation. In addition, the performance improvement of five-sequence Code MSA over three is not markedly significant. Therefore, constructing a three-sequence Code MSA presents a faster alternative while maintaining comparable accuracy. This efficiency makes it a viable option in situations where speed is essential, presenting a practical balance between effectiveness and efficiency.
Ablation Study of Token Semantic Enhancer. To validate the effectiveness of Token Semantic Enhancer, we conduct an ablation study assessing the contributions of its two components: the token-type projection layer and the token-type aware attention mechanism. We evaluate three configurations: the complete framework, a variant with only the attention module, and a baseline with both components removed. Results in Table 6 indicate that the full model achieves the highest F1 score. Removing the projection layer while retaining attention leads to a dramatic performance drop to $37.8\%$, worse than disabling both components ($85.4\%$). Attention alone performs $47.6\%$ worse than removing both projection and attention, implying that unguided attention washes out type distinctions. This further underscores the effectiveness of token-type projection layers in capturing type-specific semantic patterns. In addition, the performance of removing both components illustrates that even with Codeformer capturing token correlations, the lack of token-type information severely constrains performance, underscoring the need for both projection layers and attention for type-aware semantics. The experimental results confirm that both components are essential and must operate jointly to enable effective type-aware semantic modeling.
Evaluation of the Similarity and Optimization Module. We further examine the Classification and Optimization module by evaluating combinations of similarity measures and loss functions. Specifically, we consider three widely used similarity measures: Late Interaction, Cosine Similarity, and Euclidean Distance. In addition, we incorporate two loss functions, namely Margin Loss and Binary Cross-Entropy (BCE) Loss. Table 7 shows that the combination of Late Interaction and Margin Loss achieves the highest F1 score. Cosine Similarity performs notably worse, likely due to its insensitivity to magnitude and distributional shifts that are critical for distinguishing fine-grained semantic differences. It’s worth noting that Euclidean Distance paired with Margin Loss also offers competitive results. Considering its lower computational overhead relative to Late Interaction, Euclidean Distance offers an efficiency–effectiveness trade-off, rendering it particularly suitable for time-critical scenarios.
Table 6. Ablation study of Token Semantic Enhancer
Method Precision Recall F1
Token-Type Projection + Attention (AlphaCC) 97.8 96.3 97.1
Only Attention 55.4 28.7 37.8
Without Projection and Attention 85.8 84.9 85.4
Table 7. Different Combinations of Similarity Measures and Loss Functions
Similarity Measure Loss Function Precision Recall F1
Late Interaction Margin Loss 97.8 96.3 97.1
Cosine Similarity Margin Loss 75.1 97.2 84.7
Euclidean Distance Margin Loss 97.3 92.4 94.8
Late Interaction BCE Loss 95.6 95.1 95.3
Cosine Similarity BCE Loss 66.4 93.1 77.5
Euclidean Distance BCE Loss 94.8 87.6 91.1
5.5. RQ4: Robustness and Generalization
To evaluate AlphaCC’s robustness and generalization capability, which is critical for real-world applicability, we perform cross-dataset validation using three representative datasets: GCJ, BigCloneBench, and OJClone. Inspired by prior research practices (Xu et al., 2024), the model is trained on a source dataset and fine-tuned on a small subset ($10\%$) of the target dataset before testing. As reported in Table 8, AlphaCC consistently maintains high performance across all transfer settings. For example, when trained on OJClone and evaluated on GCJ, the model achieves an F1 score of $96.3\%$, maintaining top-tier performance relative to all baselines. These results confirm that AlphaCC does not overfit to dataset-specific characteristics. Instead, it learns generalizable semantic representations that transfer effectively across diverse code distributions, demonstrating strong robustness and practical utility in real-world software environments.
Table 8. Robustness analysis on AlphaCC
Dataset Performance of AlphaCC
Training set Test set Precision Recall F1
OJClone GCJ 98.3 94.3 96.3
BigCloneBench GCJ 94.3 92.3 93.3
GCJ OJClone 96.3 91.8 94
BigCloneBench OJClone 97.4 93 95.2
GCJ BigCloneBench 88.3 94.2 91.2
OJClone BigCloneBench 89.6 91.8 90.7
6. DISCUSSION
6.1. Interpretation and Implication
In Section 3, we briefly outline our motivations for developing AlphaCC and explain our biological analogy-based intuition for its effectiveness. Here, we provide a more comprehensive interpretation of the factors contributing to AlphaCC’s superior performance. These insights may serve as foundational principles for future research aiming to develop even more powerful code clone detection tools.
(a) Recursive Implementation
(b) Iterative Implementation
Figure 6. Two Implementations of GCD
[…]
GCD example
Multiple Sequence Alignment Mechanism. To enhance the semantic features of token sequences, AlphaCC introduces the innovative concept of “Code MSA”, which retrieves lexically similar code fragments from a large codebase to align with the original token sequence. This approach significantly elevates the model’s capacity to capture semantic information embedded in code.
Specifically, as illustrated in Figure 6, when presented with two code implementations for computing the greatest common divisor (GCD), one recursive and one iterative, the MSA mechanism can identify that these are semantic clones despite their structural differences. The critical operations in GCD functions revolve around the modulo operation a%b and the base condition check if(b==0). By incorporating similar code fragments that share these key operational characteristics but differ in other aspects, the MSA mechanism creates a rich learning environment where the model can distinguish between essential and non-essential code patterns. This data augmentation process, with its introduction of relevant variations, trains the model to focus on functionally significant code segments while ignoring irrelevant syntactic variations, thereby enhancing its sensitivity to semantically equivalent code despite structural dissimilarities.
Multiple Attention Mechanism. The CodeFormer employs a sophisticated dual-attention mechanism that operates at multiple levels of abstraction. The inner-sequence self-attention mechanism captures the relational dependencies between tokens within each individual token sequence, modeling control flow and data flow patterns. Concurrently, the inter-sequence cross-attention mechanism establishes connections across aligned positions in different sequences, enabling the model to identify patterns of token-level variations that preserve semantic equivalence.
This hierarchical attention framework allows AlphaCC to simultaneously reason about local code structures and their global functional implications (Wang et al., 2016; Zadeh et al., 2018). For example, when processing different loop constructs that accomplish the same task, the inner-sequence attention captures the loop’s internal logic, while the cross-attention recognizes the functional equivalence across different implementation styles. This multi-dimensional understanding of code significantly enhances the model’s ability to detect semantically equivalent code fragments regardless of their syntactic representation.
Streamlined Target of Token. AlphaCC leverages robust semantic modeling while maintaining superior computational efficiency and multi-language applicability compared to more complex approaches, such as AST-based or graph-based methods (Yu et al., 2019; Zhang et al., 2019; Xu et al., 2024; Yu et al., 2023). By directly analyzing code tokens, AlphaCC eliminates the need for language-specific analyzers or intricate structural transformations. This streamlined approach offers strong adaptability, enabling effective application to diverse and complex scenarios, such as multi-language software repositories or domain-specific programming environments.
6.2. Threats To Validity
Despite our rigorous evaluation of AlphaCC, several threats to validity arise due to practical limitations in resource accessibility.
First, our comparative analysis faces limitations in including certain baseline methods due to insufficient implementation details in original publications. For example, CC2Vec (Dou et al., 2024), a token-based approach, could not be accurately replicated because the training methodology for token type weights is not explicitly described in the available literature. Similarly, DeepSim (Zhao and Huang, 2018), a graph-based method, lacks comprehensive details about its feature matrix transformation process, which hinders our ability to implement it faithfully. These information gaps prevent us from incorporating these potentially relevant methods as baselines in our experimental setup in a consistent way, which may affect the comprehensiveness of our comparative evaluation.
Second, our evaluation of AlphaCC relies predominantly on three widely recognized datasets: GCJ, BigCloneBench, and OJClone. Although these datasets represent standard benchmarks in code clone detection research, they may not fully capture the diverse spectrum of code instances encountered in real-world software development environments. The characteristics of these datasets might introduce biases that influence performance metrics. Furthermore, to be more specific, although AlphaCC is designed as a language-agnostic framework operating on tokenized source code, these benchmark datasets only contain limited programming language types. In future work, we plan to extend the evaluation to a broader spectrum of programming languages to fully validate AlphaCC’s multi-language applicability claims.
7. CONCLUSION
In this paper, we propose AlphaCC, a novel approach designed to tackle the challenges of limited capability for semantic code clone detection and strong reliance on language-specific analyzers. AlphaCC first transforms each input code fragment into a token sequence. To enrich token semantic representation, AlphaCC then constructs Code MSA to align lexically similar code fragments, thereby exposing their functional commonalities. The aligned token sequences are then processed by a dual-attention Codeformer module, which leverages self-attention to capture token-level dependencies within each individual token sequence, and cross-attention to model interactions across multiple aligned sequences, enabling a robust abstraction of underlying program logic. Finally, AlphaCC computes fine-grained token-level similarity between code fragments to effectively distinguish clone pairs. Comprehensive experiments on three diverse datasets demonstrate the multi-language applicability and superior semantic understanding of AlphaCC. Furthermore, AlphaCC achieves competitive efficiency, underscoring its practical utility for large-scale code clone detection tasks.